In 2000, Amy Holland started a small mountaineering gear retail store in Northern Colorado. The store rapidly became known as Holland Gear. Ten years later, Amy had opened four additional retail stores that we covering the major regions in the US. The main product line was a family of in-house designed high-tech jackets that we fully waterproof while being breathable and lightweight. These versatile jackets sold throughout the entire year and their demand was not materially affected by the time of the year. The jackets were highly valuable products that were account for approximately 78% of total sales. Annual jackets revenues for each store were of similar size and amounted to roughly $1,000,000 per store.

After some years of operation, Holland Gear had decided to focus on its core competencies of design and retail and to outsource production to a textile contract manufacturer in Northern Mexico. Holland jackets sold at an average retail price of $325, which represented a mark-up of 30% above what Holland paid the manufacturer.

The current supply chain operated as follows. Each store made its own inventory decisions. Stores were supplied directly from the contract manufacturer by truck. The manufacturer as part of long-term contract that not only stipulated a variable unit sourcing cost, but also transportation characteristics offered the following transportation package. A shipment of up to a full truckload, which was more than 3000 jackets, was charged a flat fee of $2,200 and was guaranteed to reach the store in two weeks after order placement. (While the stores were located at slightly different distances from the manufacturer, a uniform average flat fee was agreed upon to offer uniform service to each store.) Typically, stores placed roughly two orders per year, each of about 1500 jackets.

Looking into the future, Amy was contemplating a new initiative for Holland Gear. The plan would be to transform the brick-and-mortar retail business into an Internet store. This would involve closing down the five retail stores and opening up one fulfillment center to serve the entire market. Amy felt that consolidated inventory management should provide considerable savings over the current setup.

[Adopted from Palu Gear case study by J. A. Van Mieghem, Kellogg Business School]

  1. Determine the following parameters of the inventory management system used by Holland Gear per store:

Demand rate (per year):

Demand rate (per week):

Holding cost per unit (the holding cost percentage is 20% of the value):

Fixed ordering cost per order:

Lead time (in weeks):

Current order quantity:

Current time between orders:

Reorder point:

Total number of stores:

For the Questions 2 and 3 assume that demand is steady and predictable.

  1. What is the current inventory policy used by Holland at each store? Calculate the total cost of the current inventory policy for each store and the total.
  1. Is there a better order size? Calculate the total cost of an inventory policy with the optimal order size for each store and the total. Does the total cost lower or higher than the total cost from Question 2?
  1. In reality, demand fluctuates from week to week. In fact, past weekly demand at each store exhibited a standard deviation of about 30 jackets. How should the inventory policy be adjusted if each Holland store wants its non-stockout probability be equal to 95%? Would you recommend to use Q or P review system? (Explain). Calculate the total cost of a new inventory policy for each store and the total. Assume the backorder cost is $75 per unit.

  1. What could be advantages and disadvantages from adopting the Internet initiative?
Answer #1
Demand Rate (per year) 3000
Demand Rate (per week ) 60
Holding cost per unit 6
Fixed ordering costs per order 2200
Lead time (in weeks) 2
Current order quantity 1500
Current time between orders 6 months
Re-order point =average weekly usage * lead time
=60*2
=120/week
Total number of stores 10
Cost Price Quantity 1500
Variable costs 250 375000
Holding costs 50 75000
Ordering costs 2200 2200
452200
Cost Price Quantity
Variable costs 250 3000 750000
Holding costs 50 3000 150000
Ordering costs 2200 2 4400
904400
Holding costs = 20% of COGS
=20% of 250
=50

Better order quantity

EOQ= (2 * annualdemand * costperorder) Holdingcost

EOQ = (2 * 3000 * 2200) 50

EOQ = (13200000) 50

EOQ = V264000

EOQ =513.80=514

Cost Price Quantity 1500 Quantity 514
Variable costs 250 375000 128500
Holding costs 50 75000 25700
Ordering costs 2200 4400 13200
454400 167400
Holding costs = 20% of COGS
=20% of 250
=50
No of orders =3000/1500 =3000/514
2 5.836575875
Odering costs 2200 =2200*2 =2200*6
4400 13200

The total costs at better quantity are far lower than current order level costs

P Model in inventory control is also known as fixed period system That means orders are placed at fixed interval on time. It could be once in days, weeks or years
Once order interval is fixed, order quantity is to be decided which depends on rate of demand/consumption and lead time for replanishment
Q model in inventory, the order is placed whenever the inventory level reaches a certain level, where inventory is regulary monitered, the order is trigerred whenever inventory reaches a re-order pint
The company wants to maintain non-stockout and hence Q model is recommended
Numer of weeks between order 2
Square root N 1.41
Standard deviation of weekly demand 30
Probability demand of non-stockout 0.95
Standard normal 60
Area under the curve of 95% using appendix 1 1.65
Safety stock Square root of N* Non-stockout probability* standard deviation
=1.41*30*1.65
=69.795
Reorder quantity =(average weekly demand)*(lead time in weeks)+ SS
=(57.69*2)+Safety stock
=115.38+69.795
=185.175
Q model
Cost Price Quantity 185
Variable costs 250 46250
Holding costs 50 9250
Ordering costs 2200 37400
92900
No of orders =3000/185
16.21621622